# Adjoint log canonical foliated singularities on surfaces

### v1

@@ 4.1. Outline of the classification
The classification proceeds by a systematic analysis of the intersection numbers $K\cdot C$ for irreducible components $C\subset E$.
We first study components satisfying $K\cdot C\leq 0$. Such curves are shown to be extremely restricted: they are either invariant or non-invariant curves of specific numerical types, which are completely classified in Lemmas 4.2 and 4.3. These curves form the basic building blocks of the exceptional divisor.
Starting from components with $K\cdot C<0$, we then introduce certain linear configurations, called special $K^{>0}$-chains. These chains play a central role in the analysis. A key feature is their rigidity: only finitely many numerical types can occur, and these are explicitly classified in Proposition 4.5.
Once all special chains are identified, we subtract their contribution from $K$ and consider the modified divisor
$K_{1}:=K-\sum_{F}M(K,F),$
where the sum runs over all special $K^{>0}$-chains $F\subset E$. To simplify the analysis and to avoid interactions between different chains, we impose a technical assumption ()(*), under which all special $K^{>0}$-chains are pairwise disjoint. (The case where two special chains overlap is treated separately in Proposition 4.10.) Under this assumption, the remaining components of $E$ can be analyzed according to the sign of $K_{1}\cdot C$.
We then classify the irreducible components that are not contained in any special $K^{>0}$-chain according to whether they are $\mathcal{F}$-invariant and to the sign of $K_{1}\cdot C$. This leads to strong numerical restrictions on their self-intersection numbers and valencies in the dual graph. Combining these results yields a complete list of possible configurations for $E$ under assumption ()(*), stated in Proposition 4.13.
Finally, we show that this list is exhaustive for $\epsilon$-adjoint log canonical singularities. This leads to a finite classification of exceptional divisors, summarized in Theorem 4.14.
@@ 4.2. $K$-non-positive curves
Let $C$ be an irreducible curve contained in $E$. We define the contributions coming from non-invariant and invariant components by
@@
Moreover, $KC\leq 0$ if and only if $C$ belongs to one of the following cases:
@@
      $C$ is an elliptic Gorenstein leaf. In this case, $KC=0$.
@@
By Proposition 2.8,
$K_{\mathcal{F}}\cdot C=2p_{a}(C)-2+{\rm Z}(\mathcal{F},C).$
Thus,
@@
where ${\rm Z}(\mathcal{F},C)-d_{2}(C)\geq 0$.
If $p_{a}(C)\geq 1$, then $KC\geq 0$. Moreover, $KC=0$ if and only if $p_{a}(C)=1$ and $K_{\mathcal{F}}\cdot C=d(C)=0$. In this case, $C$ is an elliptic Gorenstein leaf. This is case (1-a).
Next we assume $p_{a}(C)=0$. Then
@@
and $KC\leq 0$ implies $d(C)\leq 2$.
  • If $d(C)=0$, then $d_{2}(C)=0$ and
@@
    Since $C^{2}<0$ and $p_{a}(C)=0$, we have ${\rm Z}(\mathcal{F},C)\geq 1$, by Proposition 2.8 and Lemma 2.9.
    • If ${\rm Z}(\mathcal{F},C)\geq 3$, then $KC\geq 1-2\epsilon>0$.
    • If ${\rm Z}(\mathcal{F},C)=2$, then $C$ is a (2)(-2)-$\mathcal{F}$-curve and $KC=-2\epsilon$. This is case (1-b).
    • If ${\rm Z}(\mathcal{F},C)=1$, then $C$ is a (1)(-1)-$\mathcal{F}$-curve and $KC=-(1+2\epsilon)$. This is case (1-c).
  • If $d(C)=1$, then
@@
    • If ${\rm Z}(\mathcal{F},C)-d_{2}(C)\geq 2$, then $KC\geq 1-\epsilon>0$.
    • If $d_{2}(C)=0$ and ${\rm Z}(\mathcal{F},C)=1$, then $C$ is a (1)(-1)-$\mathcal{F}$-curve with $KC=-\epsilon$. This is case (2-b).
    • If $d_{2}(C)=1$ and ${\rm Z}(\mathcal{F},C)=1$, then $C$ is a (1)(-1)-$\mathcal{F}$-curve with $KC=-(1+\epsilon)$. This is case (2-a).
    • If $d_{2}(C)=1$ and ${\rm Z}(\mathcal{F},C)=2$, then $C$ is a (2)(-2)-$\mathcal{F}$-curve with $KC=-\epsilon$. This is case (2-c).
  • If $d(C)=2$, then $KC={\rm Z}(\mathcal{F},C)-d_{2}(C)\geq 0$. Moreover, $KC=0$ if and only if ${\rm Z}(\mathcal{F},C)=d_{2}(C)\in\{1,2\}$.
    • If ${\rm Z}(\mathcal{F},C)=d_{2}(C)=1$, then $C$ is a (1)(-1)-$\mathcal{F}$-curve. This is case (3-a).
    • If ${\rm Z}(\mathcal{F},C)=d_{2}(C)=2$, then $C$ is a (2)(-2)-$\mathcal{F}$-curve. This is case (3-b).
We now turn to the non-$\mathcal{F}$-invariant case.
Lemma 4.3.
(4.4) $KC={\rm tang}(\mathcal{F},C)+\epsilon(2p_{a}(C)-2)+d_{1}(C)+\epsilon d(C).$
Moreover, $KC\leq 0$ if and only if ${\rm tang}(\mathcal{F},C)=0$, and $C$ is one of the following explicitly listed curves:
@@
    $d(C)=0$.
    1. (a)
      $C$ is a smooth elliptic curve. In this case, $KC=0$.
    2. (b)
      $C$ is a smooth rational curve. In this case, $KC=-2\epsilon$.
@@
We compute
$KC$ $=(K_{\mathcal{F}}+E^{\rm n\!-\!inv})\cdot C+\epsilon(K_{X}+E)\cdot C$
$KC={\rm tang}(\mathcal{F},C)+\epsilon\bigl(2p_{a}(C)-2\bigr)+d_{1}(C)+\epsilon d(C).$
$KC=(d(C)-2)\epsilon.$
Then $KC\leq 0$ implies case (1-b), (2) or (3). ∎
@@ 4.3. Special $K^{>0}$-chains in $E$
Let $C$ be an irreducible component of $E$. By Lemmas 4.2 and 4.3, any irreducible component $C\subset E$ with $KC<0$ satisfies $d(C)\leq 1$. If $d(C)=0$, then necessarily $E=C$. Otherwise, when $d(C)=1$, such a curve intersects ECE-C in a unique point, and hence naturally gives rise to a linear configuration inside $E$. This motivates the introduction of the following notion.
@@
Suppose $E$ has at least two components. Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{>0}$-chain with first curve $C_{1}$. In all cases below, we have
$d(C_{1})=1,\qquad d(C_{i})=2\ \text{and}\ KC_{i}=0\ \text{for all }i\geq 2.$
Moreover, $F$ belongs to one of the following configurations:
@@
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve, and $C_{i}$ is a (2)(-2)-$\mathcal{F}$-curve for all $i\geq 2$.
@@
    There exists some $k\in[1,r]$ such that $C_{k}$ is a non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{k})=0$. Moreover, $C_{k-1}$ (if $k\geq 2$) and $C_{k+1}$ (if $k\leq r-1$) are (1)(-1)-$\mathcal{F}$-curves, and all remaining $C_{i}$ are (2)(-2)-$\mathcal{F}$-curves.
In addition,
@@
Since $KC_{1}<0$ and $d(C_{1})\geq 1$, by Lemmas 4.2 and 4.3, the curve $C_{1}$ must fall into one of the following cases:
@@
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$. In this case, $KC_{1}=-(1+\epsilon)$.
@@
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=1$ and $d_{2}=0$. In this case, $KC_{1}=-\epsilon$.
@@
    $C_{1}$ is a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$. In this case, $KC_{1}=-\epsilon$.
@@
    $C_{1}$ is a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{1})=0$, $d_{1}=0$ and $d_{2}=1$. In this case, $KC_{1}=-\epsilon$.
If $r=1$, then $F=C_{1}$ and the statement follows directly from the above list. Hence we assume $r\geq 2$.
Since $KC_{i}\geq 0$ for all $i\geq 2$, Proposition 3.8 implies
$KC_{k}<\frac{-KC_{1}}{\mu_{k}},\qquad k\geq 2,$
where $\mu_{k}=\det\big(-C_{i}C_{j}\big)_{1\leq i,j\leq k-1}$.
Cases (ii)–(iv). In these cases $KC_{1}=-\epsilon$, hence $0\leq KC_{i}<\epsilon$ for all $i\geq 2$. By (4.3) and (4.4), for each $i\geq 2$ we have $KC_{i}=0$, and $C_{i}$ must be one of the following:
@@
A simple combinatorial analysis yields the types (B), (C), and (D), and any additional case is excluded by the separatrix theorem.
Case (i). Here $KC_{1}=-(1+\epsilon)$ and $d(C_{1})=d_{2}(C_{1})=1$. In particular, we have $C_{1}^{2}\leq-2$, and hence
$0\leq KC_{k}<\frac{1+\epsilon}{2}<1-\epsilon,\qquad k\geq 2.$
Since $d_{2}(C_{1})=1$, the curve $C_{2}$ is $\mathcal{F}$-invariant. Moreover, as $0\leq KC_{2}<1-\epsilon$ and $d_{2}(C_{2})\geq 1$, it follows from (4.3) that $d(C_{2})=d_{2}(C_{2})=2$ and $KC_{2}=0$. Thus $C_{2}$ is either a (2)(-2)-$\mathcal{F}$-curve or a (1)(-1)-$\mathcal{F}$-curve. The latter is excluded by the separatrix theorem.
By the same argument, each curve $C_{i}$ with $i\geq 3$ must be a (2)(-2)-$\mathcal{F}$-curve satisfying $d(C_{i})=d_{2}(C_{i})=2$ and $KC_{i}=0$. Therefore, the chain $F$ is of type (A). ∎
@@
Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{\geq 0}$-chain in $E$ with $KC_{1}<0$. If $C$ is an irreducible curve in $E\setminus F$ such that $F\cdot C=C_{r}\cdot C=1$, then
@@
4.4. Classification of the exceptional divisor
@@ 4.4.1. Case I: $E$ irreducible
If $G_{E}=\{[C]\}$, then $C$ belongs to one of the following cases:
@@
It follows directly from Lemma 4.2 and Lemma 4.3. ∎

### v2

@@ 4.1. Outline of the classification
The classification is based on the analysis of the intersection numbers $K\cdot C$ for irreducible components $C\subset E$.
Conceptually, the classification can be viewed as a finite assembly problem. Once all irreducible components with prescribed intersection behaviour are classified, the remaining configurations arise from combining a fixed finite collection of admissible building blocks, subject to the numerical constraints imposed by negative definiteness and the adjoint inequalities.
We first classify all components with $K\cdot C\leq 0$, which are shown to be highly restricted and are completely described in Lemmas 4.2 and 4.3. As a consequence, two elementary cases can be separated and treated independently, namely the irreducible case and the $K^{=0}$-graph case (Propositions 4.9 and 4.10). After excluding these cases, we may assume that there exists a component $C\subset E$ with $K\cdot C<0$.
Starting from such components, we introduce special linear configurations, called special $K^{>0}$-chains, whose possible numerical types are finite and classified in Proposition 4.6. Subtracting their contribution from $K$, we obtain a modified divisor $K_{1}$, and—under a disjointness assumption ()(*)—reduce the analysis to the sign of $K_{1}\cdot C$ for the remaining components. The case where ()(*) fails is treated separately in Proposition 4.11.
Using Lemmas 4.12 and 4.13, the case $K_{1}\cdot C<0$ is reduced to explicitly described configurations. Hence, it suffices to analyze the case where $E$ is a $(K_{1})^{=0}$-graph, which reduces to a finite combinatorial problem. This yields the complete list of possible configurations (Proposition 4.14), and hence the classification of $\epsilon$-adjoint log canonical singularities (Theorem 4.15).
@@ 4.2. $K$-non-positive curves
Let $C$ be an irreducible curve contained in $E$. We define the following intersection-theoretic quantities:
@@
We then analyze complements satisfying $K\cdot C\leq 0$. We first consider the non-$\mathcal{F}$-invariant case.
Lemma 4.2.
(4.3) $KC=\epsilon\bigl(2p_{a}(C)-2+d(C)\bigr)+{\rm tang}(\mathcal{F},C)+d_{1}(C).$
Moreover, the inequality $KC\leq 0$ forces ${\rm tang}(\mathcal{F},C)=0$, and under this condition $C$ is necessarily one of the following explicitly listed types:
  1. (1)
    $d(C)=0$.
    1. (a)
      $C$ is a smooth elliptic curve. In this case, $KC=0$.
    2. (b)
      $C$ is a smooth rational curve. In this case, $KC=-2\epsilon$.
  2. (2)
  3. (3)
Proof.
We compute
$KC$ $=(K_{\mathcal{F}}+E^{\rm n\!-\!inv})\cdot C+\epsilon(K_{X}+E)\cdot C$
$KC={\rm tang}(\mathcal{F},C)+\epsilon\bigl(2p_{a}(C)-2+d(C)\bigr)+d_{1}(C).$
$KC=(d(C)-2)\epsilon.$
Then $KC\leq 0$ implies case (1-b), (2) or (3). ∎
We now turn to the $\mathcal{F}$-invariant case.
@@
Furthermore, the inequality $KC\leq 0$ holds if and only if $C$ is one of the following curves:
@@
      $C$ is an irreducible elliptic Gorenstein leaf. In this case, $KC=0$.
@@
By Proposition 2.8, we have
$K_{\mathcal{F}}\cdot C=2p_{a}(C)-2+{\rm Z}(\mathcal{F},C),$
so that
@@
with ${\rm Z}(\mathcal{F},C)-d_{2}(C)\geq 0$.
First, if $p_{a}(C)\geq 1$, then clearly $KC\geq 0$. Moreover, $KC=0$ if and only if $p_{a}(C)=1$ and ${\rm Z}(\mathcal{F},C)=d(C)=0$, in which case $C$ is an irreducible elliptic Gorenstein leaf, corresponding to case (1-a).
Now assume $p_{a}(C)=0$. Then
@@
so that $KC\leq 0$ implies $d(C)\leq 2$. Since $C^{2}<0$ and $p_{a}(C)=0$, Proposition 2.8 together with Lemma 2.9 ensures ${\rm Z}(\mathcal{F},C)\geq 1$.
Case $d(C)=0$. Here $d_{2}(C)=0$ and
@@
Given $\epsilon\in(0,1/3)$, $KC\leq 0$ forces ${\rm Z}(\mathcal{F},C)\in\{1,2\}$, corresponding to cases (1-b) and (1-c).
Case $d(C)=1$. In this case, $d_{2}(C)\in\{0,1\}$, and
@@
Since $\epsilon\in(0,1/3)$, $KC\leq 0$ implies
$({\rm Z}(\mathcal{F},C),d_{2}(C))=(1,1),\ (1,0),\ (2,1),$
which correspond to cases (2-b), (2-b) and (2-c), respectively.
Case $d(C)=2$. Then
$KC={\rm Z}(\mathcal{F},C)-d_{2}(C)\geq 0,$
and $KC=0$ occurs precisely when ${\rm Z}(\mathcal{F},C)=d_{2}(C)\in\{1,2\}$, corresponding to cases (3-a) and (3-b). ∎
Corollary 4.4.
Let $C\subset E$ be an irreducible curve with $d(C)\geq 1$.
@@
    If $0\leq KC<\epsilon$, then $KC=0$.
@@
Write
$KC=\begin{cases}m\epsilon+n,&\text{if $C$ is non-$\mathcal{F}$-invariant},\\[2.84526pt] m(1+\epsilon)+n,&\text{if $C$ is $\mathcal{F}$-invariant},\end{cases}$
with $m=2p_{a}(C)-2+d(C)\geq-1$ and $n\geq 0$, as in (4.3) and (4.4).
For non-$\mathcal{F}$-invariant $C$, we have m=1m=-1 implies $KC=-\epsilon+n\in\{-\epsilon,1-\epsilon,2-\epsilon,\dots\}$, $m=0$ implies $KC=n\in\{0,1,2,\dots\}$, and $m\geq 1$ implies $KC\geq\epsilon$.
For $\mathcal{F}$-invariant $C$, we have m=1m=-1 implies $KC=-(1+\epsilon)+n\in\{-(1+\epsilon),-\epsilon,1-\epsilon,2-\epsilon,\dots\}$, $m=0$ implies $KC=n\in\{0,1,2,\dots\}$, and $m\geq 1$ implies $KC\geq 1+\epsilon$.
Since $\epsilon\in(0,1/3)$, $0\leq KC<\epsilon$ forces $KC=0$ for non-$\mathcal{F}$-invariant $C$, and $0\leq KC<(1+\epsilon)/2$ forces $KC=0$ for $\mathcal{F}$-invariant $C$. ∎
@@ 4.3. Special $K^{>0}$-chains in $E$
Let $C$ be an irreducible component of $E$. By Lemmas 4.2 and 4.3, any irreducible component $C\subset E$ with $KC<0$ satisfies $d(C)\leq 1$. If $d(C)=0$, then necessarily $E=C$. Otherwise, when $d(C)=1$, such a curve intersects ECE-C in a unique point, and hence naturally gives rise to a linear configuration inside $E$. This motivates the introduction of the following notion. These chains capture the configurations where a single curve with $KC<0$ is followed by curves with $KC\geq 0$, forming a linear segment of $E$.
@@
Suppose $E$ has at least two components. Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{>0}$-chain with first curve $C_{1}$. Then $F$ belongs to one of the following configurations:
@@
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve, and $C_{i}$ is (2)(-2)-$\mathcal{F}$-curve for all $i\geq 2$.
@@
    There exists some $k\in[1,r]$ such that $C_{k}$ is a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{k})=0$. In this case, $C_{k-1}$ (if $k\geq 2$) and $C_{k+1}$ (if $k\leq r-1$) are (1)(-1)-$\mathcal{F}$-curves, and all remaining $C_{i}$ are (2)(-2)-$\mathcal{F}$-curves.
Moreover, in all cases above, the first curve $C_{1}$ satisfies $d(C_{1})=1$, while every other component $C_{i}$ ($i\geq 2$) satisfies
$d(C_{i})=2\quad\text{and}\quad KC_{i}=0.$
Furthermore,
@@
Since $KC_{1}<0$ and $d(C_{1})\geq 1$ (as $E$ has at least two components), Lemmas 4.2 and 4.3 imply that $C_{1}$ must fall into one of the following cases:
@@
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$, in which case $KC_{1}=-(1+\epsilon)$;
@@
    $C_{1}$ is a (1)(-1)-$\mathcal{F}$-curve with $d_{1}=1$ and $d_{2}=0$, in which case $KC_{1}=-\epsilon$;
@@
    $C_{1}$ is a (2)(-2)-$\mathcal{F}$-curve with $d_{1}=0$ and $d_{2}=1$, in which case $KC_{1}=-\epsilon$;
@@
    $C_{1}$ is a smooth rational non-$\mathcal{F}$-invariant curve with ${\rm tang}(\mathcal{F},C_{1})=0$, $d_{1}=0$, and $d_{2}=1$, in which case $KC_{1}=-\epsilon$.
If $r=1$, then $F=C_{1}$, and the statement follows directly from the above list. Hence we may assume that $r\geq 2$.
Since $F$ is a special $K^{>0}$-chain, Proposition 3.6 yields
$0\leq KC_{k}<\frac{-KC_{1}}{\mu_{k}},\qquad k\geq 2,$
where $\mu_{k}=\det\bigl(-C_{i}C_{j}\bigr)_{1\leq i,j\leq k-1}$.
Cases (ii)–(iv). In these cases $KC_{1}=-\epsilon$, hence $0\leq KC_{i}<\epsilon$ for all $i\geq 2$. By Corollary 4.4, this implies $KC_{i}=0$. Consequently, for each $i\geq 2$, the curve $C_{i}$ must be one of the following:
@@
A straightforward combinatorial analysis of the possible adjacencies then yields the types (B), (C), and (D), and any additional case is excluded by the separatrix theorem (cf. Theorem 2.10).
Case (i). In this case, $KC_{1}=-(1+\epsilon)$ and $d(C_{1})=d_{2}(C_{1})=1$. Since $C_{1}$ is not $\mathcal{F}$-exceptional, we have $\mu_{2}=-C_{1}^{2}\geq 2$. It follows that $C_{2}$ is $\mathcal{F}$-invariant, as $d(C_{1})=d_{2}(C_{1})=1$, and
$0\leq KC_{2}<\frac{1+\epsilon}{\mu_{2}}\leq\frac{1+\epsilon}{2}.$
By Corollary 4.4, this implies $KC_{2}=0$. Consequently, $C_{2}$ is either a (2)(-2)-$\mathcal{F}$-curve or a (1)(-1)-$\mathcal{F}$-curve, by Lemma 4.3. The latter possibility is excluded by the separatrix theorem. Thus, $d(C_{2})=d_{2}(C_{2})=2$. Moreover, $C_{2}^{2}\leq-2$, since otherwise $C_{2}$ would be $\mathcal{F}$-exceptional, contradicting our assumptions. Therefore, $C_{1}+C_{2}$ is an $\mathcal{F}$-chain, and in particular $\mu_{3}\geq 2$.
Applying the same argument to $C_{3}$, we see that $C_{3}$ is $\mathcal{F}$-invariant, $d(C_{3})=d_{2}(C_{3})=2$, and $KC_{3}=0$, so that $C_{1}+C_{2}+C_{3}$ is again an $\mathcal{F}$-chain. Iterating this procedure, we conclude that $F$ is an $\mathcal{F}$-chain with the first curve $C_{1}$, hence of type (A). ∎
@@
Let $F=C_{1}+\cdots+C_{r}$ be a special $K^{\geq 0}$-chain in $E$ with the first curve $C$. If $C$ is an irreducible curve in $E\setminus F$ such that $F\cdot C>0$, then $F\cdot C=C_{r}\cdot C=1$ and
@@
Proof.
The statement is a direct consequence of Proposition 3.6 and Proposition 4.6, including the arguments in its proof. ∎
4.4. Classification
@@ 4.4.1. Case I: $E$ irreducible
If $E=C$ is irreducible, then $C$ belongs to one of the following cases:
@@
This corresponds to the case $d(C)=0$. The classification follows directly from Lemmas 4.2 and 4.3. ∎